KyledaGreat:

Graph the original function by indicating how the more basic function has been shifted, reflected, stretched, or compressed.

KyledaGreat:

1 attachment
KyledaGreat:

1 attachment
KyledaGreat:

1 attachment
KyledaGreat:

@extrinix @vocaloid

Extrinix:

Okay so you understand that the quadratic equation looks like this \(f(x)=\sqrt{x-h}+k\) With \(h\) being the opposite of it's seen value, and \(k\) being it's seen value In this equation it shows this \(s(x)=-\sqrt{x+4}-5\) So, you can see that \(h\) is \(4\), meaning that it's value on the graph is \(-4\), meaning it `shifts left 4` As well as that, you can see that \(k\) is \(-5\), meaning that the graph `shifts down 5` Now, there is a negative next to the square root, meaning that the graph is `flipped across the x-axis` Y-axis flips would only occur if \(s(x)\) was negative, meaning that there is `no change` in it's y-axis placement

KyledaGreat:

Okay , what about vertical stretch/compress and the vertical shift ?

KyledaGreat:

@extrinix

Extrinix:

As shown, the equation \(f(x)=a\sqrt{x-h}+k\) \(a\) represents the stretch or compression of the equation, but as \(a\) is equal to 1 on your equation, there is `no change` in it's stretch or compression

KyledaGreat:

okay good

KyledaGreat:

1 attachment
KyledaGreat:

1 attachment
Extrinix:

👍👍

KyledaGreat:

Consider the following function. \[u(x) =-\sqrt{x+4}\] Determine the domain and range of the original function. Express your answer in interval notation.

Extrinix:

Use my previous post as an example and try it for yourself